Let f(x) = x + 2, g(x) = (x/(x+5)), and h(x) = x^2 – 5x – 3
Find: (f + g)(x)

Answer by stanbon(75887)

Let f(x) = x + 2, g(x) = (x/(x+5)), and h(x) = x^2 – 5x – 3
Find: (f + g)(x)
--------
(f+g)(x) = f(x)+g(x) = (x+2) + (x/x+5) = [(x+5)(x+2) + x]/(x+5)
----
= [x^2 + 8x + 10]/(x+5)
-----------
Cheers,
Stan H.
------------